### v3

# Sample size and power calculations for causal inference in observational studies

Abstract
This paper investigates the theoretical foundation and develops analytical formulas for sample size and power calculations for causal inference with observational data. By analyzing the variance of the inverse probability weighting estimator of the average treatment effect, we decompose the power calculations into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that to determine the minimal sample size of an observational study, it is sufficient under mild conditions to have two parameters additional to the standard inputs in the power calculation of randomized trials, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose using the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. Utilizing the Lyapunov Central Limit Theorem on the linear combination of covariates, our procedure does not require distributional assumptions on the multivariate covariates. We develop an associated R package PSpower.
Keywords: Causal inference; Observational study; Overlap; Power; Sample size; Weighting
1 Introduction
Randomized controlled trials are the gold standard for causal inference. Designing randomized trials requires sample size and power calculations, in which investigators pre-specify a target estimand—typically the effect size of a treatment—and characteristics of the sample of the intended study. Then the goal is to determine the minimal sample size needed to achieve the statistical power ($1-\beta$), given a type I error rate ($\alpha$) level, in testing the treatment effect estimated by a specific method based on a sample that satisfies these characteristics, or conversely, to determine the power in detecting a treatment effect at the $\alpha$ level given a fixed sample size. Sample size and power calculations are the two sides of the same coin and we will use them interchangeably henceforth. Power calculation is conducted in the design stage of a study, prior to data collection. Thus, individual data are usually unavailable; investigators at best have some summary information of previous similar studies, e.g. mean and variances of covariates and outcomes by treatment arm. Analysts must rely on limited summary information of the intended study to recover the full data structure. The power calculation for randomized trials with treatment at one time has been well established and is routinely performed (Chow et al.,, 2017; Branson et al.,, 2024).
When randomized trials are not feasible due to logistical or ethical constraints, researchers are increasingly using observational data to emulate randomized trials. The power analysis for observational studies has been less developed in the literature. In fact, its role and utility has been questioned by some authors (Hernán,, 2022), who argued that the primary consideration in observational studies is to reduce confounding bias rather than improve efficiency or power. However, we view power analysis as crucial in designing prospective observational studies, which are usually subject to cost and time constraints. For example, in studies of rare diseases, recruiting patients is time-consuming, and thus investigators must determine the minimal sample size to achieve a reliable estimate before conducting the study. Another example concerns commercial population health data (such as the Truveta electronic health record data or IBM’s MarketScan health insurance claims data), which usually charge users cost proportional to the data size. With limited resources, a sample size calculation is critical for designing observational studies based on such data sources. Power analysis also helps researchers to understand the scope of data required and guide database selection in the design stage.
Power analysis for causal inference with observational studies is challenging because of confounding. In randomized trials, there is no confounding by design; therefore, one can simply use the difference in the mean outcomes between arms to estimate the treatment effects. Closed-form formulas, e.g., based on the two-sample z-test, are readily applicable for sample size calculation. In contrast, in observational studies, confounding must be adjusted to obtain an unbiased estimate of treatment effects. Applying power calculation formulas designed for randomized trials to observational studies often leads to severely underpowered studies. In practice, a general power analysis method is to use simulations to generate individual data based on pre-specified summary statistics (Qin,, 2024). This method often makes unverifiable distributional assumptions of the multivariate covariates and lacks theoretical basis on what summary statistics are adequate to simulate the entire data. In fact, many different data generating models may satisfy the same summary statistics but lead to vastly different sample sizes; choosing between these models is usually ad hoc. Another method is to use the variance inflation factor (Hsieh and Lavori,, 2000; Scosyrev and Glimm,, 2019; Shook-Sa and Hudgens,, 2022; Austin,, 2021) to approximate the potential inflation in the sample size of an observational study compared to a randomized trial. In particular, Shook-Sa and Hudgens, (2022) derived a rigorous variance inflation factor formula based on the analysis of the variance of the inverse probability weighting estimator of the average treatment effect; their formula requires the distribution of the weights, which is often not available.
In summary, the existing power calculation methods usually require information on the data generating process or the distributions of the weights or covariates, which is not always available in the design stage of a study. In this paper, we bypass this limitation by–under mild conditions–reducing the variance formula to a function of a few interpretable and commonly available summary parameters instead of the weights or covariates. Based on the analysis of variance of the inverse probability weighting estimator, we decompose power calculations into three components: propensity score distribution, potential outcome distribution, and their correlation. We show that, to determine the minimal sample size of an emulated trial with treatment at a single time, in addition to the standard inputs in the power calculation of a randomized trial, it is sufficient to have two parameters, which quantify the strength of the confounder-treatment and the confounder-outcome association, respectively. For the former, we propose to use the Bhattacharyya coefficient, which measures the covariate overlap and, together with the treatment proportion, leads to a uniquely identifiable and easily computable propensity score distribution. For the latter, we propose a sensitivity parameter bounded by the R-squared statistic of the regression of the outcome on covariates. We appeal to the Lyapunov Central Limit Theorem to reduce the multivariate covariates into a single Gaussian variable, which allows our procedure not to require specific distributional assumptions of the covariates. We show that a special case of our formula is the formula by Shook-Sa and Hudgens, (2022), which formalized the confounder-treatment association but not the confounder-outcome association. We develop an associated R package PSpower, available at https://CRAN.R-project.org/package=PSpower.
2 Variance and sample size based on the Hájek estimator
Suppose we have a sample of $N$ units. Each unit $i$ ($i=1,2,\ldots,N$) has a binary treatment indicator $Z_{i}$, with $Z_{i}=0$ being the control and $Z_{i}=1$ being treated, and a vector of $p$ covariates $X_{i}=(X_{1i},\cdots,X_{pi})^{\prime}$. For each unit $i$, there are two potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$ mapped to treatment and control status, of which only the one corresponding to the observed treatment is observed, denoted by $Y_{i}=Z_{i}Y_{i}(1)+(1-Z_{i})Y_{i}(0)$; the other potential outcome is missing. The propensity score is the probability of a unit being assigned to the treatment group given the covariates (Rosenbaum and Rubin,, 1983): $e(x)=\Pr(Z_{i}=1\mid X_{i}=x)$.
The goal is to use the sample to estimate the average treatment effect:
$\tau^{\mbox{\tiny{ATE}}}=\operatorname{\mathbb{E}}\{Y_{i}(1)-Y_{i}(0)\}=\operatorname{\mathbb{E}}_{x}\{\tau(x)\},$ (1)
where $\tau(x)=\operatorname{\mathbb{E}}\{Y_{i}(1)-Y_{i}(0)\mid X_{i}=x\}$ is the conditional average treatment effect at covariate value $x$. To identify $\tau^{\mbox{\tiny{ATE}}}$, we maintain the two standard assumptions: (i) unconfoundededness, $\Pr(Z_{i}\mid Y_{i}(0),Y_{i}(1),X_{i})=\Pr(Z_{i}\mid X_{i})$; (ii) overlap, i.e. $10<\Pr(Z_{i}\mid X_{i})<$ 1. Then we can identify the $\tau^{\mbox{\tiny{ATE}}}$ from the observed data as $\tau^{\mbox{\tiny{ATE}}}=\operatorname{\mathbb{E}}\{Y_{i}Z_{i}/e(X_{i})\}-\operatorname{\mathbb{E}}\{Y_{i}(1-Z_{i})/(1-e(X_{i}))\}$. Among the consistent estimators of $\tau^{\mbox{\tiny{ATE}}}$, we focus on the Hájek inverse probability weighting estimator for analytical simplicity:
$\widehat{\tau}_{N}=\frac{\sum_{i=1}^{N}Y_{i}Z_{i}/{e}(X_{i})}{\sum_{i=1}^{N}Z_{i}/{e}(X_{i})}-\frac{\sum_{i=1}^{N}Y_{i}(1-Z_{i})/\{1-{e}(X_{i})\}}{\sum_{i=1}^{N}(1-Z_{i})/\{1-{e}(X_{i})\}}.$ (2)
Lunceford and Davidian, (2004) show that $N^{1/2}(\widehat{\tau}_{N}-\tau^{\mbox{\tiny{ATE}}})$ is asymptotically normal with variance
$V=\mathbb{E}\left[\frac{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{[Y(0)-\mathbb{E}\{Y(0)\}]^{2}}{1-e(X)}\right].$ (3)
In observational studies, the true propensity score is unknown and is replaced by an estimated one. Lunceford and Davidian, (2004) show that the asymptotic variance of the Hájek estimator with estimated propensity score, $V_{0}$, is $V$ subtracted by a complex quadratic term and therefore smaller. Calculation of the quadratic term usually requires individual data, which is unavailable in most sample size calculation. Therefore, we will proceed with $V$ henceforth, noting that it leads to a conservative (i.e. larger) sample size estimate than that based on $V_{0}$. We will investigate the discrepancy between $V$ and $V_{0}$ using simulations in Section 4.
With a fixed value of $V$, a $(1-\alpha)$ confidence interval for $\widehat{\tau}_{N}$ is $\widehat{\tau}_{N}\pm z_{1-\alpha/2}V^{1/2}N^{-1/2}$, where $z_{1-\alpha/2}$ is the $(1-\alpha/2)$ quantile of the standard normal distribution. Without loss of generality, we consider an one-sided level-$\alpha$ test for the null hypothesis $H_{0}:\tau=0$ against the alternative $H_{a}:\tau>0$:
reject $H_{0}$ if $\widehat{\tau}_{N}>z_{1-\alpha}({V}/N)^{1/2}$. (4)
Standard derivation shows the minimal sample size $N$ to reach a desired power $\beta$ for a level-$\alpha$ test (4) with a fixed variance $V$ is (the proof is relegated to Supplementary Material A.1):
$N={\tau^{-2}}{\left(z_{1-\alpha}{V}^{1/2}-z_{1-\beta}{V_{0}}^{1/2}\right)^{2}}\leq{\tau^{-2}}{V}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (5)
In a randomized experiment, the propensity score is fixed and known, equaling the treatment proportion $r$, so that ${V}$ simplifies to $\left[\operatorname{\mathbb{V}}\{Y(1)\}+\operatorname{\mathbb{V}}\{Y(0)\}\right]/\{2r(1-r)\}$. Moreover, $\Pr\{Y(z)\}=\Pr(Y\mid Z=z)$, and thus variance of the potential outcomes $Y(z)$ can be estimated by the sample variances of the outcome in arm $z$, $s_{z}^{2}$. Following the convention, we impose $s_{0}=s_{1}=s$ and define the effect size as $\widetilde{\tau}=\tau/s$. Then, to detect an effect size $\widetilde{\tau}$ in a randomized experiment using a one-sided level-$\alpha$ test (4) with power $\beta$, the minimal sample size is
$N=\frac{s_{1}^{2}+s_{0}^{2}}{r(1-r)\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}=\frac{\left(z_{1-\alpha}+z_{\beta}\right)^{2}}{r(1-r)\widetilde{\tau}^{2}}.$ (6)
This coincides with the standard sample size formula of a two-sample $z$-test.
In observational studies, power calculation is more complicated because (i) the propensity scores $e(X)$ are not known, (ii) $\Pr(Y(z))\neq\Pr(Y\mid Z=z)$ due to confounding, and thus the terms involving potential outcomes, $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$, deviate from their observed counterparts $\operatorname{\mathbb{E}}(Y\mid Z=z)$ and $\operatorname{\mathbb{V}}(Y\mid Z=z)$, and (iii) the numerator and denominator inside the expectation operator in (3) can not be separated because the potential outcomes may be dependent of the propensity scores. When sampled units are available, we can readily address these complications. However, power calculation is conducted prior to data collection, with little information to estimate the propensity scores or calculate the expectations. Nonetheless, analysts usually have some summary information from historical studies about the intended population.
3 Sample size and power calculation
The essence of sample size calculation lies in computing the variance $V$ in (3). If we could observe $(X_{i},Z_{i},Y_{i})$ for each sampled unit $i$, $V$ could easily be estimated by sample moments, but not so with only summary data. To determine which summary information is sufficient for computing $V$, we rewrite the terms in expression (3) as (taking $z=1$ for example),
$\mathbb{E}\left[\frac{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}{e(X)}\right]=\mathbb{E}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}}\right]\mathbb{E}\left\{\frac{1}{e(X)}\right\}+\mathrm{Cov}\left[{[Y(1)-\mathbb{E}\{Y(1)\}]^{2}},\frac{1}{e(X)}\right].$ (7)
This expression suggests that we need three components: (i) mean and variance of the potential outcome in each arm, (ii) mean of (the inverse of) the propensity score, and (iii) covariance between the potential outcomes and propensity score. We study each component in detail below.
3.1 Normal approximation of linear combinations of covariates
Both the propensity score and the outcome models are commonly specified as a generalized linear model with a linear combination of the covariates, $X^{\prime}\beta=\sum_{j=1}^{p}\beta_{j}X_{j}$ as the predictor. However, in power calculation, we usually do not have individual data or even summary statistics of each covariate, without which we cannot characterize the distributions of the propensity score or the outcome. Previous authors often operated with a single normal predictor (e.g. Raudenbush,, 1997; Kelcey et al.,, 2017; Qin,, 2024), but without theoretical justification. Here we present a special case of the Lyapunov Central Limit Theorem (Billingsley,, 1995) to approximate the distribution of the scalar summary $X^{\prime}\beta$ and thus justify such a procedure.
Theorem 1.
Assume $X_{1},X_{2},\dots$ is a sequence of independent random variables with mean 0 and variance 1. Let $\beta_{1},\beta_{2},\dots$ be a sequence of real numbers. If the following two conditions hold: (i) there exists a positive number $B$ such that $\mathbb{E}(X_{j}^{4})<B$, and (ii) $\max_{1\leq j\leq p}{\beta_{j}^{2}}(\sum_{j=1}^{p}\beta_{j}^{2})^{-1}=o(1)$, then
$\frac{1}{\sum_{j=1}^{p}\beta_{j}^{2}}\sum_{j=1}^{p}\beta_{j}X_{j}\Rightarrow\mathsf{N}(0,1).$ (8)
The proof of Theorem 1 is relegated to Supplementary Material A.2. Condition (i) ensures that the distribution of each covariate does not have heavy tails, and condition (ii) ensures that no single covariate dominates the linear combination. Though these two conditions may not always hold, empirical evidence from simulated and three real applications (see Supplementary Material B.1) shows the normality generally holds in a wide range of scenarios.
3.2 Distribution of propensity scores
The propensity score is usually specified by a logistic regression model: $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}=\beta_{0}+\sum_{j=1}^{p}\beta_{j}X_{j}$. Theorem 1 motivates us to approximate the distribution of the scalar summary $W_{e}$ by a normal distribution $\mathsf{N}(\mu_{e},\sigma_{e}^{2})$, and thus approximate the propensity score distribution by a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$.
To determine the two parameters $\mu_{e}$ and $\sigma_{e}$ of the logit-normal distribution, we would need to impose constraints in the form of summary information of the treatment assignment. The first constraint, which is usually available, is the proportion of the treatment group in the intended study: $r=\Pr(Z_{i}=1)$. The second constraint is a measure, denoted as $\phi$, of the similarity between the covariate distributions in the treatment and control groups, $\Pr(X\mid Z=1)$ versus $\Pr(X\mid Z=0)$. There are many choices of the specification of $\phi$; important selection criteria include: (i) there should be 1-1 map between $(r,\phi)$ and $(\mu_{e},\sigma_{e})$ so that the data generating process of the propensity score is unique, and (ii) easy to compute the propensity score distribution given $(r,\phi)$.
The logit-normal distribution does not have a closed-form mean and thus it is generally difficult to solve for $(\mu_{e},\sigma_{e})$ given $(r,\phi)$. Instead, we propose to approximate the logit-normal distribution by a Beta distribution, which has a range on $[0,1]$ and is easy to compute. Applying the method in Aitchison and Shen, (1980), we can derive the “optimal” Beta function approximation of a logit-normal function by minimizing the Kullback-Leibler divergence between them. Specifically, for a logit-normal distribution $\mathcal{P}(\mu_{e},\sigma_{e}^{2})$ and its optimal approximated Beta distribution $\mathsf{Beta}(a,b)$, their parameters must satisfy the following relation: $\mu_{e}=\psi(a)-\psi(b)$ and $\sigma_{e}^{2}=\psi^{\prime}(a)+\psi^{\prime}(b)$, where $\psi$ and $\psi^{\prime}$ are the digamma and trigamma functions, respectively. Numerical examples in Supplementary Material B.2 illustrate that this approximation is accurate for a wide range of logit-normal distributions.
If $e(X)\sim\mathsf{Beta}(a,b)$, then there is a closed-form relation between $a,b,r$: $r={a}/(a+b)$. Therefore, it suffices to identify the propensity score distribution if we find another easily computable relation between $a,b$ and the overlap measure $\phi$. The question now becomes how to specify such a $\phi$. Directly operating on the multivariate covariate $X$ is challenging; instead we operate on the propensity score $e(X)$, which is the coarsest balancing score (Rosenbaum and Rubin,, 1983). Let $f_{z}(u)$ be the density of the propensity score in arm $z$: $f_{z}(u)=\Pr(e(X)=u\mid Z=z)$. We propose to use the Bhattacharyya coefficient (Bhattacharyya,, 1943) to measure the covariate overlap between two groups:
$\phi\equiv\int_{0}^{1}\{f_{1}(u)f_{0}(u)\}^{1/2}\,\mathrm{d}u.$ (9)
The Bhattacharyya coefficient $\phi$ satisfies $0\leq\phi\leq 1$, with $\phi=0$ if and only if the support of $f_{0}$ and $f_{1}$ does not overlap, whereas $\phi=1$ if and only if $f_{0}=f_{1}$ almost everywhere. Therefore, $\phi$ is an appropriate measure of overlap; we term it as the overlap coefficient hereafter. We choose the Bhattacharyya coefficient because it has a major computable advantage in our context, established in the following proposition (the proof is relegated to Supplementary Material A.3).
Proposition 1.
For $e(X)\sim\mathsf{Beta}(a,b)$, the overlap coefficient $\phi$ has a closed-form expression as a function of $(a,b)$:
$\phi=\frac{\Gamma(a+0.5)}{{a}^{1/2}\Gamma(a)}\frac{\Gamma(b+0.5)}{{b}^{1/2}\Gamma(b)}.$ (10)
Therefore, given $(r,\phi)$, we can obtain the Beta parameters $(a,b)$ by solving for Equations $r={a}/(a+b)$ and (10). Specifically, based on the former, we can reparameterize $(a,b)$ as $a=kr$ and $b=k(1-r)$, under which Equation (10) becomes
$\phi(k)=\frac{\Gamma(kr+0.5)}{(kr)^{1/2}\Gamma(kr)}\frac{\Gamma(k(1-r)+0.5)}{\{k(1-r)\}^{1/2}\Gamma(k(1-r))}.$ (11)
We write $\phi(k)$ to highlight that, given $r$, the overlap coefficient $\phi$ is a function of the unknown $k$. The following proposition establishes that, given $r$, there is a 1-1 map between $\phi$ and $k$, and equivalently, a 1-1 map between $(r,k)$ and $(a,b)$. The proof is relegated to Supplementary Material A.4.
Proposition 2.
The function, $\phi(k)$, defined in (11), is monotonically increasing in $k$ if $k>\max\left\{(2r)^{-1},\{2(1-r)\}^{-1}\right\}$.
Proposition 2 allows us to use the bisection method to solve for $k$ efficiently. Interestingly, we found that many other specification of the overlap coefficient do not lead to a 1-1 map to the propensity score distribution. In summary, as long as users specify $(r,\phi)$, we can solve for the Beta parameters $(a,b)$ in a computationally efficient fashion, based on which we further reverse the Beta approximation of $e(X)$ to obtain the logit-normal distribution parameters $(\mu_{e},\sigma_{e})$.
To visualize the overlap coefficient, Figure 1 shows the distributions of $e(X)$ versus $1-e(X)$ corresponding to different combinations of $(r,\phi)$. In practice, we recommend users to combine substantive knowledge with Figure 1 to provide a reasonable input of $\phi$ given $r$. For example, as a generic rule of thumb for a balanced design with $r=0.5$, we recommend to code $\phi=(0.9,0.8,0.5)$ as good, medium, and poor overlap, respectively.
Figure 1: Distributions of $e(X)$ versus $1-e(X)$ corresponding to different combinations of $(r,\phi)$, where $e(X)$ follows a $\mathsf{Beta}(a,b)$ distribution with $a$ and $b$ determined by $(r,\phi)$.
3.3 Distribution of potential outcomes
For simplicity, we focus on continuous outcomes. The potential outcomes are usually specified as a linear model $Y(z)=\gamma_{z0}+X^{\prime}\gamma_{z}+\epsilon_{z}$, with $\mathbb{E}[\epsilon_{z}]=0$. Similarly as before, Theorem 1 motivates us to approximate the distribution of $X^{\prime}\gamma_{z}$ by a Gaussian random variable $W_{z}$.
Recall that we approximated the distribution of the propensity score by $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ is a normal random variable. Therefore, the correlation between the potential outcomes and the propensity score is captured by the correlation between $W_{e}$ and $W_{z}$. This motivates us to decompose $W_{z}$ into two parts, one linear in $W_{e}$ and the other orthogonal to $W_{e}$. Specifically, we can decompose $W_{z}=a_{z}W_{e}+W_{e}^{\perp}$ such that $\mathrm{cov}(W_{e},W_{e}^{\perp})=0$. Then, the potential outcomes can be expressed as $Y(z)=a_{z}W_{e}+W_{e}^{\perp}+\epsilon_{z}$, where both $W_{e}^{\perp}$ and $\epsilon_{z}$ are unknown normal random variables. Because both variables are independent of $W_{e}$, it is impossible to further identify $W_{e}^{\perp}$ from $\epsilon_{z}$; thus, with slight abuse of notation, we combine these two terms and denote the sum as $\epsilon_{z}$, under which the potential outcomes are modeled as $Y(z)=a_{z}W_{e}+\epsilon_{z},\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma^{2}_{z}).$
Here we have six unknown parameters in total to characterize the potential outcome distributions: $a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2}$. In sample size calculation settings, users usually have prior knowledge of the mean and variance of the observed outcome in each treatment group (e.g. available in Table 1 of a previous study in medical papers), denoted as $E_{z}\equiv\mathbb{E}(Y\mid Z=z)$ and $S_{z}^{2}\equiv\operatorname{\mathbb{V}}(Y\mid Z=z)$, for $z=0,1$. These are different from their counterparts to the potential outcomes $\operatorname{\mathbb{E}}\{Y(z)\}$ and $\operatorname{\mathbb{V}}\{Y(z)\}$. In addition to these four quantities, we would need two more constraints to identify the six parameters; it is natural to use the correlation between the covariates and the observed outcomes, i.e., $\rho_{z}\equiv\mathrm{cor}(Y,W_{e}\mid Z=z)$. We postpone the detailed discussion of $\rho_{z}$ to Section 3.4.
With the six specified values $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, one can solve for the unknown parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ by the following equations (the proof is relegated to Supplementary Material A.5)
$a_{z}$ $=\rho_{z}|S_{z}|\{\operatorname{\mathbb{V}}(W_{e}\mid Z=z)\}^{-1},\quad\mu_{z}=E_{z}-a_{z}\mathbb{E}(W_{e}\mid Z=z),\quad\sigma_{z}^{2}=(1-\rho_{z}^{2}){S_{z}^{2}}.$ (12)
The conditional mean $\mathbb{E}(W_{e}\mid Z=z)$ and variance $\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ do not have closed forms. We propose to calculate them numerically by the following formula
$\mathbb{E}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x},$ (13)
$\operatorname{\mathbb{V}}(W_{e}\mid Z=z)$ $=\frac{\int_{\mathbb{R}}x^{2}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}{\int_{\mathbb{R}}\exp\left\{-(2\sigma_{e}^{2})^{-1}(x-\mu_{e})^{2}\right\}\operatorname{\mathrm{e}xpit}\{(-1)^{1-z}x\}\,\mathrm{d}x}-\left\{\mathbb{E}(W_{e}\mid Z=z)\right\}^{2}.$ (14)
Once the parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ are solved, we can fully determine the distribution of $Y(0)$ and $Y(1)$.
3.4 Correlation between outcomes and covariates
We now interpret the seemingly opaque correlation parameter $\rho_{z}=\mathrm{cor}(Y,W_{e}\mid Z=z)$. Recall that the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, thus $W_{e}$ summarizes the covariates that are predictive of the treatment assignment. Therefore, $\rho_{z}$ measures the confounder-outcome association: as $\rho_{z}$ increases, confounding covariates exert a greater influence on the outcome between treatments. The parameter $\rho_{z}$ is intrinsically related to a standard and interpretable quantity: the covariate-outcome association, $R_{z}=\mathrm{cor}(Y,\widetilde{Y}\mid Z=z)$, where $\widetilde{Y}$ is the component of $Y$ that can be explained by the covariates. Under the linear outcome model $Y(z)=X^{\prime}\gamma_{z}+\epsilon_{z}$, we have $\widetilde{Y}=X^{\prime}\gamma_{z}$ and the square of the covariate-outcome association, $R_{z}^{2}$, is equivalent to the R-squared statistic $R_{z,Y\mid X}=\operatorname{\mathbb{V}}(X^{\prime}\gamma_{z}\mid Z=z)/\operatorname{\mathbb{V}}(Y\mid Z=z)$. Then, we can derive the following expression to show that $\rho_{z}$ is bounded by $R_{z}$:
$\rho_{z}^{2}$ $=\left[\frac{\mathrm{cov}(Y,W_{e}\mid Z=z)}{\{\mathbb{V}(W_{e}\mid Z=z)\}^{1/2}\{\mathbb{V}(\widetilde{Y}\mid Z=z)\}^{1/2}}\right]^{2}{\frac{\mathbb{V}(\widetilde{Y}\mid Z=z)}{\mathbb{V}(Y\mid Z=z)}}=\mathrm{cor}^{2}(\widetilde{Y},W_{e}\mid Z=z)R_{z}^{2}\leq R^{2}_{z}.$ (15)
In randomized trials, covariates are independent of the treatment assignment by design and thus $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)=0$ and $\rho_{z}=0$. In observational studies, as the covariates associated with propensity scores differ more from the covariates associated with potential outcomes, $\mathrm{cor}(\widetilde{Y},W_{e}\mid Z=z)$ will get closer to 0. Equation (15) implies that users can bound $\rho_{z}^{2}$ by the $R_{z}^{2}$ by the R-squared value of a hypothetical regression of the outcome on the covariates, of which investigators often have an educated guess based on prior domain knowledge. Moreover, $\rho_{z}$ is invariant under linear transformation of the potential outcome models. Specifically, under homogeneous effects $Y(1)=Y(0)+\tau$, we always have $\rho_{1}=\rho_{0}$; under heterogeneous effects, simulations and real examples in Section 4 suggest that $\rho_{1}$ and $\rho_{0}$ are usually similar. Also, as will be shown in formula (17), the minimal sample size for $\tau^{\mbox{\tiny{ATE}}}$ depends on $\rho$ only through $\rho^{2}$. Therefore, in practice, we recommend to impose $\rho_{1}=\rho_{0}=\rho$ for simplicity, and conduct a sensitivity analysis of $|\rho|$ in the range $(0,|R_{Y|X}|)$ with a domain knowledge informed specification of $R_{Y|X}$, or set $\rho=0$ when such information is unavailable, which could lead to a mild underestimate of the sample size.
3.5 Analytical formula for sample size calculation
Summarizing the building blocks in Section 3.2-3.4, we propose to approximate the distribution of the propensity score and the potential outcomes as
$e(X)=\mathrm{expit}(W_{e}),\quad Y(z)=a_{z}W_{e}+\epsilon_{z},\quad W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2}),\quad\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2}),$ (16)
respectively, where the propensity score model parameters $(\mu_{e},\sigma_{e})$ and the outcome model parameters $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$ can be uniquely determined by the user specified inputs $(r,\phi)$ and $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$, respectively. To further reduce the number of inputs, we impose $S_{1}^{2}=S_{0}^{2}=S^{2}$ and $\rho_{1}=\rho_{0}=\rho$. Empirical evidence from multiple real data, e.g. the ones investigated in the supplementary materials, supported this restriction. Furthermore, following the convention in previous power analysis, we define the standardized effect size as the treatment effect scaled by the standard error: $\widetilde{\tau}=\tau/S$.
The following theorem presents the analytical formula for the sample size calculation of the $\tau^{\mbox{\tiny{ATE}}}$ (the proof is relegated to Supplementary Material A.6).
Theorem 2.
Let $\widetilde{\tau}$ be the desired standardized effect size of the ATE. The minimum sample size for a two-sided level-$\alpha$ test (4) for detecting $\widetilde{\tau}$ with power $\beta$ is $N=\widetilde{V}\left(z_{1-\alpha/2}+z_{\beta}\right)^{2}/{\widetilde{\tau}^{2}}$, with
$\widetilde{V}=$ $\rho^{2}\left(\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}\right)+2(1-\rho^{2})+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}+(1-\rho^{2})\right\}\exp(-\mu_{e}+\sigma_{e}^{2}/2)$
+$+\left\{\rho^{2}\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}+(1-\rho^{2})\right\}\exp(\mu_{e}+\sigma_{e}^{2}/2),$ (17)
where $\sigma_{e\mid z}^{2}=\mathbb{V}(W_{e}\mid Z=z)$.
Theorem 2 shows that the formula of $\widetilde{V}$ does not depend on $\mu_{z}$. Moreover, equation (12) shows that $E_{z}$ is involved in solving $\mu_{z}$ but not $a_{z}$ and $\sigma_{z}^{2}$. Therefore, calculating $\widetilde{V}$ does not require $E_{z}$, further reducing the number of inputs. Indeed, the formula of $\widetilde{V}$ only involves $\sigma_{e}^{2}$, $\mu_{e}$ and $\sigma_{e\mid z}^{2}$, all of which can be derived from the distribution of $e(X)$ that is determined by $(r,\phi)$, and an additional input $\rho$. This simplification reduces the number of inputs for the sample size calculation to two, namely $(\phi,\rho)$, in addition to the standard $\widetilde{\tau},\alpha,\beta,r$. In the special case of a randomized trial, $\rho=0$, and the minimal sample size reduces to formula (6).
Remark 1. The above analytical derivation elucidates the connection and distinction between sample size calculations in randomized trials and observational studies. Specifically, four parameters — $\alpha$, $\beta$, the standardized effect size $\widetilde{\tau}$, and the treatment proportion $r$—are required in both cases. For observational studies, two additional inputs are required: the overlap coefficient $\phi$ and the correlation parameter $\rho^{2}$, both of which have natural interpretation. This result is intuitive because the key difference between observational studies and randomized trials is the presence of confounders in the former, which are pre-treatment covariates that affect the treatment assignment and the outcome simultaneously. The parameters $\phi$ and $\rho^{2}$ quantify the strength of confounder-treatment and confounder-outcome correlation, respectively. Similar strategy is used in the formulation of the E-value (Ding and VanderWeele,, 2016). Intuitively, as the overlap decreases (i.e., $\phi$ deceases) and the correlation between covariates and the outcome increases (i.e., $\rho^{2}$ increases), the observational study deviates further from a randomized trial, rendering the marginal distribution of the observed outcomes deviate further from that of the potential outcomes and thus require more sample units to recover the causal estimands. Indeed, the calculated sample size monotonically increases with $\rho^{2}$ from (17) (the proof is relegated to Supplementary Material A.7.). Moreover, numerical experiments show that the calculated sample size monotonically decreases with $\phi$.
Remark 2. The sample size formula of Shook-Sa and Hudgens, (2022) – derived based on the variance inflation factor – is a special case of formula (17). Specifically, when assuming the propensity score distribution $e(X)\sim\operatorname{\mathrm{e}xpit}(\mathsf{N}(\mu_{e},\sigma_{e}^{2}))$, their formula becomes $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$, which is formula (17) with $\rho=0$. We show that variance $\widetilde{V}_{\mathrm{SH}}$ is smaller than the variance $\widetilde{V}$ in (17) by a term proportional to $\rho^{2}$ (details in Supplementary Material A.7). Although $\rho^{2}$ is small in many of our simulation studies, ignoring it risks underestimating the sample size in general settings. Therefore, we recommend to report power calculations with a range of $\rho^{2}$ values between 0 and $R^{2}_{Y|X}$. Shook-Sa and Hudgens, (2022) discussed a remainder term that implicitly adjusts for the confounder-treatment association, but did not point out the connection, and the remainder term is only estimable with prior data on both weights and outcomes. In contrast, our method does not require this information and explicitly formulates the role of the confounder-treatment association.
Remark 3. For binary outcomes, because the mean of a Bernoulli distribution determines the variance, compared to the continuous outcomes in Section 3.3, we have two fewer constraints to determine the outcome distribution. As a simple solution, we recommend to use linear models for binary outcomes so that the proposed method directly applies. Numerical examples in Section 4.2 and Supplementary Material D.3 suggest that this approach generally works well in practice.
4 Numerical Examples
4.1 Synthetic Data
Following the simulation design in Li et al., (2019), we simulate $N=1,000,000$ units with 10 covariates of different types as the superpopulation. Covariates $X_{1}$ to $X_{4}$ are binary with $\mathsf{Ber}(p)$ with $p=0.2,0.4,0.6,0.8$ respectively; $X_{5}\sim\mathsf{Unif}(0,1)$; $X_{6}$ to $X_{8}$ follows a Poisson distribution with mean parameters $1,2,$ and $3$ respectively; $X_{9}\sim\mathsf{Gamma}(2,3)$ and $X_{10}\sim\mathsf{Beta}(2,3)$. We generate the treatment indicator $Z\sim\mathsf{Ber}(\operatorname{\mathrm{e}xpit}(\beta_{0}+\kappa X\beta))$, where $\beta=(1,1,-1,0,-2,1,0.5,0,0,0)^{\prime}$ and $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ is the parameter that controls the overlap between two treatment groups. For these $\alpha$ values, we choose $\beta_{0}=\{0,-0.248,-0.489,-0.722,-0.860,-0.951\}$ accordingly so that the treatment proportion $r=\mathbb{E}(Z)=0.5$. We then simulate the potential outcome $Y(z)\sim\mathsf{N}(X\gamma+\tau z,4^{2})$, where $\gamma=(1,1,-1,-1,0,-1,-1,0,1,1)^{\prime}$, and let the observed outcome $Y_{i}=Z_{i}Y(1)+(1-Z_{i})Y(0)$. The outcome model implies a homogeneous treatment effect $\tau(x)=1$ and thus all average effects including $\tau^{\mbox{\tiny{ATE}}}$ equal 1.
The overlap between two treatment groups is shown in Figure 2: $\kappa=0$ corresponds to randomized trials; as $\kappa$ increases, the overlap decreases. The values of $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$ correspond to the overlap coefficient $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, respectively.
Figure 2: Distribution of the propensity scores in two treatment arms with fixed $r=0.5$ and various overlap coefficient $\phi=1.00,0.98,0.93,0.87,0.84,0.81$, respectively.
Now we fix the power at 0.80, and solve for the sample size under each value of $\kappa$. We calculate the eight summary statistics $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ as discussed in Section 3 from the $N=1,000,000$ units in the superpopulation. The proportion of treatment, $r$, is calculated by the sample mean $\overline{Z}$, and the observed mean and variance, $E_{1},E_{0},S_{1}^{2},S_{0}^{2}$ are calculated with values of $Y$ on subsets $Z=1,0$, respectively. With the logistic model for the propensity score, $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X\beta)$, the latent variable $W_{e}$ is estimated by $\widehat{\beta}_{0}+X\widehat{\beta}$, where $(\widehat{\beta}_{0},\widehat{\beta})$ is the estimated intercept and coefficients of the logistic regression of $Z$ on $X$. Although none of the covariates $X$ is normal, the latent linear combination of $X$, $W_{e}$, appears to be close to normal (see the rightmost column of Figure 4 in Supplementary Material B.1), and the correlation parameters $\rho_{z}$ can be estimated by the sample correlation between estimated $W_{e}$ and $Y(z)$. We also calculate the pooled variance $S^{2}$ and pooled correlation $\rho$ empirically. To calculate $\phi$, we use the empirical cumulative probability function of the propensity scores (details are relegated to Supplementary Material B.3). We present $(r,\phi)$ in Table 1 below, and the remaining quantities $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ and the pooled quantities $(S^{2},\rho^{2})$ are shown in Table 2 in Supplementary Material C.2. We also calculate and show in Table 2 the R-squared statistic $R^{2}$ of linear model $Y\sim X$, with $Y$ centered within the respective treatment arm. The results show that $S_{1}^{2}\approx S_{0}^{2}\approx S^{2}$ and $\rho_{1}^{2}\approx\rho_{0}^{2}\approx\rho^{2}$ in all settings. The $R^{2}$ turns out to be much larger than $\rho^{2}$, which is not surprising as $R^{2}$ is the theoretical upper bound of $\rho^{2}$, and provides the most conservative estimate when there is no substantive knowledge on the range of $\rho^{2}$.
Proposed $z$-test Shook-sa & Hudgens
$\phi$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 645 0.83 0.88 629 0.80 0.87 629 0.80 0.87
0.98 656 0.79 0.87 628 0.78 0.85 653 0.80 0.88
0.93 754 0.79 0.87 626 0.73 0.80 742 0.78 0.86
0.88 993 0.78 0.85 624 0.60 0.65 950 0.77 0.84
0.84 1290 0.80 0.85 623 0.48 0.53 1216 0.77 0.84
0.81 1617 0.81 0.87 622 0.40 0.45 1481 0.79 0.81
Table 1: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens,, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004.
We use inputs $(r,\phi,\rho^{2})$ and standardized effect size $\widetilde{\tau}=1/S$ to inversely calculate the sample size $N_{\mathrm{pos}}$. Table 1 shows the calculated sample size to achieve the nominal power 0.80 under various degrees of overlap. Then, we sample $N_{\mathrm{pos}}$ units randomly from the $N=1,000,000$ units in the superpopulation. We repeat the process for $B=10,000$ times; each time, we conduct the hypothesis test (4) at level $\alpha=0.05$, with $\widehat{\tau}_{N}$ calculated with true and estimated propensity scores, respectively. We then estimate the power by the proportion of the $B=10,000$ replications where the null hypothesis is rejected. Table 1 also reports the power calculated based on Shook-Sa and Hudgens, (2022). The proposed method consistently leads to powers close to the nominal 0.80; the method of Shook-Sa and Hudgens, (2022) leads to similar power. As noted before, the power based on the estimated propensity scores is higher than that based on the true scores, rendering the proposed method conservative. Note that the discrepancy in variance between the true and estimated propensity scores could be magnified in the simulation compared to most real studies, because both $Z$ and $Y$ are drawn from the underlying true model. We also include the sample size calculated by two-sample $z$-test for comparison. When the overlap is perfect, the $z$-test correctly calculates the sample size. However, when the overlap is poor, meaning that the intended observational study is far from a target randomized trial, the $z$-test, which treats observational studies as randomized trials and thus ignore the degree of overlap, tends to vastly underestimate the sample size.
The parameter $\rho^{2}$ is usually unknown; we suggest treating $\rho^{2}$ as a sensitivity parameter between 00 and $R^{2}$. For $\phi\in\{1.00,0.98,0.93,0.88,0.84,0.81\}$, sensitivity analyses will yield sample sizes of $(629,630)$, $(653,669)$, $(742,821)$, $(959,1239)$, $(1225,1818)$ and $(1515,2516)$, respectively. The results show that when the overlap is good, i.e., the observational study resembles a randomized trial, the minimal sample size is robust to the choice of $\rho^{2}$. However, as the overlap deteriorates, the sample size becomes more sensitive to the choice of $\rho^{2}$. This finding restates the importance of covariate overlap in observational studies, and presents the challenge of sample size calculation with worse overlap. The wide intervals do not reflect a drawback of the proposed method; instead, it displays the sensitivity to the “hidden” parameter $\rho^{2}$, which has been mostly ignored in the existing literature. The sensitivity also shows the importance of the bound $\rho^{2}\leq R^{2}$, because $R^{2}$ is often small (e.g. 0.2) in clinical studies, which guarantees that the sample size falls into a reasonable range with modest overlap.
We also conduct simulation studies where (i) the outcomes are binary and (ii) the estimand is the average treatment effect for the treated or the overlap population, the details of which are relegated to Supplementary Material D.3 and D.4, respectively. The proposed method is shown to correctly estimate the sample size under both settings.
4.2 Real data based simulations
This section illustrates the proposed methods using simulations based on a widely used benchmark data from the Right Heart Catheterization observational study (Connors et al.,, 1996), which was conducted to evaluate the causal effect of the diagnostic cardiological procedure of right heart catheterization (binary treatment) among hospitalized critically ill patients. The purpose of our simulation is not to re-design an existing study; instead, we want to assess the performance of our methods in a real world setting where we do not know the true data generating model, presenting a more realistic and challenging setting. The data contain 5735 patients, each with 51 covariates (20 continuous, 25 binary, 6 categorical). The outcome is the binary survival status at 30 days after admission, with $Y=1$ meaning death. In total, 2184 ($38.1\%$) of the patients received the treatment and the rest did not. The death rate is 38.0% and 30.6% in the treatment and control group, respectively.
The simulation proceeds as follows. We first calculate the summary parameters $(r,\phi,\rho^{2})$ and effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ from the real data. We take these values as the only inputs of a hypothetical power analysis, without access to the whole data. We apply the proposed method with these inputs to compute the minimal sample size needed to achieve a certain level of power in detecting the pre-fixed effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$. We then reverse the process to estimate the power under a sample of that calculated size. The closer the estimated power is to the nominal power (e.g. 0.8), the more accurate the our method is. We did not compare with the method of Shook-Sa and Hudgens, (2022) because it requires the distribution of weights, which is assumed not available in this study.
Following the above outline, from the real data we estimate the propensity scores using a logistic model with the main effects of each covariate (the distributions by group are shown in Supplementary Material C.1 (Figure 6). Based on these estimated scores, we numerically compute the true overlap coefficient $\rho$ defined in (9) to be 0.835. We also used the proposed normal approximation in Section 3.2, $W_{e}\sim\mathsf{N}(-0.701,2.189)$, and estimate the overlap coefficient to be 0.822, which is close to the true value, showing the precision of our proposed Beta approximation. The Hájek estimate of the standardized effect size $\widetilde{\tau}^{\mbox{\tiny{ATE}}}$ is 0.14. The treatment proportion is $r=0.381$ and $\rho^{2}=0.000$. With these inputs alone, we apply formula (17) to calculate the minimal sample size to achieve 0.8 power to be 3625. We then reverse the process to calculate the actual power under a sample of 3625 units. Specifically, we generate $B=1000$ replicates, each with 3625 rows randomly drawn from the original data with replacement. We calculate the Hájek estimate of $\tau^{\mbox{\tiny{ATE}}}$ and the associated $95\%$ confidence interval in each replicate. Among the 1000 replicates, 924 intervals do not cover zero, giving an estimated power of 0.924 with 95% confidence interval (0.908 – 0.940), which is larger than the nominal power of 0.8. Potential reasons for the discrepancy include: (i) we used the estimated instead of the true propensity scores, and (ii) we treated the binary outcome as Gaussian. We also computed the sample size calculated by the two-sample $z$-test, which is 1567, leading to a vastly underpowered estimate with the power of only 0.55. Given that any real study would have data attrition due to missing data and extreme weights, a conservative (overpowered) estimate of the sample size is arguably more desirable than the opposite.
Figure 3 shows the power curves of $\tau^{\mbox{\tiny{ATE}}}$ given different values of the overlap coefficient $\phi$ and the correlation $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$ while fixing other summary inputs as those calculated from the Right Heart Catheterization data. It shows that, to achieve the same power, the minimal sample size would increase as either the overlap decreases or the correlation $\rho$ increases. Both patterns are as expected from the theory. The $R^{2}$ value is 0.21, which serves as the upper bound of $\rho^{2}$.
We repeat the process for the average treatment effect for the treated and the overlap population, respectively. The results are relegated to Supplementary Material D.5.
Figure 3: Power and sample size curves of an emulated Right Heart Catheterization study with an effect size $\widetilde{\tau}=0.14$ under different overlap coefficient $\phi$ (left) and the correlation $\rho=\rho_{1}=\rho_{0}$ (right).
5 Discussion
This paper investigates the theoretical foundation for power analysis in the propensity score analysis of causal inference with observational data. We show that, compared with the power calculation of randomized trials, users only need to specify two parameters for emulated trials, which quantifies the strength of the confounder-treatment and the confounder-outcome association, respectively. We propose interpretable and easily computable statistics to characterize these two parameters. Our method bypasses distributional assumptions on the multivariate covariates and generalizes the existing variance inflation factor method. We provide new theoretical insights on the sufficient summary information for power calculation in observational studies.
There are several important extensions of our method in observational studies, including repeated outcomes (Liu and Wu,, 2005), time-to-event outcomes (Hsieh and Lavori,, 2000), and multilevel data (Arpino and Mealli,, 2011). Propensity weighting estimators for causal effects are available in all these settings. Therefore, we can readily adapt the general structure of power calculation developed in this paper. For example, for studies with repeated outcomes, a common estimand is the time-averaged difference; we can analyze the variance of an inverse probability weighted marginal outcome model within-subject correlation between outcomes. For time-to-event outcomes, a common estimand is the marginal causal hazard ratio from the Cox proportional hazard model; we can analyze the variance of the inverse probability weighted Cox model.
Supplementary Materials
A Proof of Propositions and Theorems
A.1 Proof of (5)
Rewrite $\widehat{\tau}_{N}=\tau+\epsilon\sqrt{V_{0}/N}$, where $\epsilon\sim\mathsf{N}(0,1)$. Then the power of test is given by
$\Pr\left(\tau+\epsilon\sqrt{\frac{V_{0}}{N}}>z_{1-\alpha}\sqrt{\frac{V}{N}}\right)$ $=\Pr\left(\epsilon>z_{1-\alpha}\sqrt{\frac{V}{V_{0}}}-\tau\sqrt{\frac{N}{V_{0}}}\right)$ (A.1)
$=1-\Phi\left(\frac{z_{1-\alpha}\sqrt{V}-\tau\sqrt{N}}{\sqrt{V_{0}}}\right).$ (A.2)
Equalizing this power with $\beta$, we obtain the sample size formula
$N=\frac{\left(z_{1-\alpha}\sqrt{V}-z_{1-\beta}\sqrt{V_{0}}\right)^{2}}{\tau^{2}}.$ (A.3)
Furthermore, when $\alpha<0.5<\beta$, $z_{1-\alpha}>0$, $z_{1-\beta}<0$, and hence we have
$N\leq\frac{V}{\tau^{2}}\left(z_{1-\alpha}-z_{1-\beta}\right)^{2}=\frac{V}{\tau^{2}}\left(z_{1-\alpha}+z_{\beta}\right)^{2}.$ (A.4)
A.2 Special case of the Lyapunov Central Limit Theorem
For completeness, we first present the general form of the Lyapunov Central Limit Theorem (CLT).
Theorem 3 (Lyapunov CLT).
Suppose $\{Z_{1},\dots,Z_{n},\dots\}$ is a sequence of independent random variables, each with finite expected value $\mu_{i}$ and variance $\sigma_{i}^{2}$. Define $s_{n}^{2}=\sum_{i=1}^{n}\sigma_{i}^{2}$. If for some $\delta>0$, the Lyapunov’s condition
$\lim_{n\rightarrow\infty}\frac{1}{s_{n}^{2+\delta}}\sum_{i=1}^{n}\mathbb{E}\left[|Z_{i}-\mu_{i}|^{2+\delta}\right]=0$ (A.5)
is satisfied, then
$\frac{1}{s_{n}}\sum_{i=1}^{n}(Z_{i}-\mu_{i})\Rightarrow\mathsf{N}(0,1).$ (A.6)
As a special case, let $Z_{i}=\beta_{i}X_{i}$, where $\{X_{1},\cdots,X_{n},\cdots\}$ is a sequence of independent random variables with expectation 0 and variance 1. Then $Z_{i}$ has expectation 0 and variance $\beta_{i}^{2}$. Assume (1) $\mathbb{E}[X_{i}^{4}]\leq B<\infty$ and (2) $M_{n}\equiv\frac{\max_{1\leq i\leq n}\beta_{i}^{2}}{\sum_{i=1}^{n}\beta_{i}^{2}}=o(1)$, then the Lyapunov’s condition as satisfied (by taking $\delta=2$) because
$\lim_{n\rightarrow\infty}\frac{1}{(\sum_{i=1}^{n}\beta_{i}^{2})^{2}}\sum_{i=1}^{n}\beta_{i}^{4}\mathbb{E}[X_{i}^{4}]$ $\leq B\lim_{n\rightarrow\infty}\frac{\sum_{i=1}^{n}\beta_{i}^{4}}{(\sum_{i=1}^{n}\beta_{i}^{2})^{2}}$
$\leq B\lim_{n\rightarrow\infty}\frac{\max_{1\leq i\leq n}\beta_{i}^{2}\sum_{i=1}^{n}\beta_{i}^{2}}{(\sum_{i=1}^{n}\beta_{i}^{2})^{2}}$
$=B\lim_{n\rightarrow\infty}M_{n}$
$=0.$
Therefore, $\sum_{i=1}^{n}\beta_{i}X_{i}$ is approximately a normal distribution with mean 0 and variance $\sum_{i=1}^{n}\beta_{i}^{2}$ when $n$ is large.
A.3 Proof of Proposition 1
Suppose $U\sim\mathsf{Beta}(a,b)$ and $Z\sim\mathsf{Bin}(U)$. Then,
$\mathbb{E}[Z]=\mathbb{E}[\mathbb{E}[Z\mid U]]=\mathbb{E}[U]=\frac{a}{a+b}.$ (A.7)
Furthermore, define $f_{z}(u)=\Pr(U=u\mid Z=z)$ for $z\in\{0,1\}$. By Bayes’ rule,
$f_{z}(u)=\Pr(U=u\mid Z=z)=\frac{\Pr(U=u)\Pr(Z=z\mid U=u)}{\Pr(Z=z)}.$ (A.8)
Therefore,
$f_{1}(u)$ $=\frac{\Pr(U=u)u}{a/(a+b)}=\frac{\Gamma(a+b)}{\Gamma(a)\Gamma(b)}\frac{u^{a}(1-u)^{b-1}}{a/(a+b)}=\frac{\Gamma(a+b+1)}{\Gamma(a+1)\Gamma(b)}u^{a}(1-u)^{b-1},$
$f_{0}(u)$ $=\frac{\Pr(U=u)(1-u)}{b/(a+b)}=\frac{\Gamma(a+b)}{\Gamma(a)\Gamma(b)}\frac{u^{a-1}(1-u)^{b}}{b/(a+b)}=\frac{\Gamma(a+b+1)}{\Gamma(a)\Gamma(b+1)}u^{a-1}(1-u)^{b}.$
Consequently,
$\phi$ $=\int_{0}^{1}\sqrt{\frac{\Gamma(a+b+1)^{2}}{\Gamma(a+1)\Gamma(b)\Gamma(a)\Gamma(b+1)}}u^{a-0.5}(1-u)^{b-0.5}\,\mathrm{d}u$
$=\frac{\Gamma(a+b+1)}{\sqrt{a}\Gamma(a)\sqrt{b}\Gamma(b)}\cdot\frac{\Gamma(a+0.5)\Gamma(b+0.5)}{\Gamma(a+b+1)}$
$=\frac{\Gamma(a+0.5)}{\sqrt{a}\Gamma(a)}\frac{\Gamma(b+0.5)}{\sqrt{b}\Gamma(b)}.$
A.4 Proof of Proposition 2
Let $h(k)=\log\phi(k)=\log\Gamma\left(kr+\frac{1}{2}\right)-\log\Gamma(kr)+\log\Gamma\left(k(1-r)+\frac{1}{2}\right)-\log\Gamma(k(1-r))-\log k-\frac{1}{2}\log(r(1-r))$. Then, it suffices to show that $h^{\prime}(k)>0$ for all $k$ such that $12kr>$ 1 and $12k(1-r)>$ 1.
Denote by $\psi$ the digamma function, $\psi(x)=(\log\Gamma(x))^{\prime}$. Then,
$h^{\prime}(k)=r\psi\left(kr+\frac{1}{2}\right)-r\psi(kr)+(1-r)\psi\left(k(1-r)+\frac{1}{2}\right)-(1-r)\psi(k(1-r))-\frac{1}{k}.$ (A.9)
Using the equality
$\psi(x)-\psi(y)=\int_{0}^{1}\frac{t^{y-1}-t^{x-1}}{1-t}\,\mathrm{d}t,\quad\forall x,y>1,$ (A.10)
we have
$h^{\prime}(k)$ $=r\int_{0}^{1}\frac{t^{kr-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t+(1-r)\int_{0}^{1}\frac{t^{k(1-r)-1}\left(1-\sqrt{t}\right)}{1-t}\,\mathrm{d}t-\frac{1}{k}$ (A.11)
$=\int_{0}^{1}\frac{rt^{kr-1}+(1-r)t^{k(1-r)-1}}{1+\sqrt{t}}\,\mathrm{d}t-\frac{1}{k}.$ (A.12)
Changing variable ($t=u^{2}$) yields
$h^{\prime}(k)$ $=\int_{0}^{1}\frac{2ru^{2kr}+2(1-r)u^{2k(1-r)}}{u(1+u)}\,\mathrm{d}u-\frac{1}{k}$ (A.13)
$=\frac{1}{k}\left\{\int_{0}^{1}\frac{\mathrm{d}\left(u^{2kr}+u^{2k(1-r)}\right)}{1+u}-1\right\}$ (A.14)
$=\frac{1}{k}\left\{\left.\frac{u^{2kr}+u^{2k(1-r)}}{1+u}\right|_{0}^{1}+\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u-1\right\}$ (A.15)
$=\frac{1}{k}\int_{0}^{1}\frac{u^{2kr}+u^{2k(1-r)}}{(1+u)^{2}}\,\mathrm{d}u>0.$ (A.16)
A.5 Proof of Equation (12).
Recall $Y(z)=a_{z}W_{e}+\epsilon_{z}$ where $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. The observed quantities satisfy the following equations.
$E_{z}$ $\equiv\mathbb{E}[Y\mid Z=z]=\mathbb{E}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}\mathbb{E}[W_{e}\mid Z=z]+\mu_{z};$ (A.17)
$S_{z}^{2}$ $\equiv\mathbb{V}[Y\mid Z=z]=\mathbb{V}[a_{z}W_{e}+\epsilon_{z}\mid Z=z]=a_{z}^{2}\mathbb{V}[W_{e}\mid Z=z]+\sigma_{z}^{2};$ (A.18)
$\rho_{z}$ $\equiv\mathrm{cor}[Y,W_{e}\mid Z=z]=\mathrm{cor}[a_{z}W_{e}+\epsilon_{z},W_{e}\mid Z=z]=\frac{a_{z}\mathbb{V}[W_{e}\mid Z=z]}{\sqrt{S_{z}\mathbb{V}[W_{e}\mid Z=z]}}.$ (A.19)
By (A.19),
$a_{z}=\rho_{z}\sqrt{\frac{S_{z}^{2}}{\mathbb{V}[W_{e}\mid Z=z]}},$ (A.20)
which, plugged into (A.17) and (A.18), yields
$\mu_{z}=E_{z}-a_{z}\mathbb{E}[W_{e}\mid Z=z],\quad\sigma_{z}^{2}=(1-\rho_{z}^{2})S_{z}^{2}.$ (A.21)
A.6 Proof of the analytical form of variance $V$ in Theorem 2.
Recall that
$V=\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]+\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right],$ (A.22)
where $e(X)=\operatorname{\mathrm{e}xpit}(W_{e})$, $Y(z)=a_{z}W_{e}+\epsilon_{z}$, $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$, $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. Therefore,
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=\mathbb{E}\left[\frac{a_{1}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{1}^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{1}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.23)
where
$\mathbb{E}\left[\frac{1}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.24)
$=\int\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.25)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.26)
$=1+\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.27)
and
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{-w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.28)
$=\int u^{2}\left(1+e^{-u-\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.29)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u+\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.30)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.31)
Therefore,
$\mathbb{E}\left[\frac{\{Y(1)-\mathbb{E}[Y(1)]\}^{2}}{e(X)}\right]=a_{1}^{2}\sigma_{e}^{2}+\sigma_{1}^{2}+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.32)
Similarly,
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=\mathbb{E}\left[\frac{a_{0}^{2}(W_{e}-\mu_{e})^{2}+\epsilon_{0}^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]=a_{0}^{2}\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (A.33)
where
$\mathbb{E}\left[\frac{1}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.34)
$=\int\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.35)
$=\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.36)
$=1+\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right),$ (A.37)
and
$\mathbb{E}\left[\frac{(W_{e}-\mu_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ $=\int(w-\mu_{e})^{2}\left(1+e^{w}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(w-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}w$ (A.38)
$=\int u^{2}\left(1+e^{u+\mu_{e}}\right)\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u$ (A.39)
$=\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{u^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u+\int u^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\sigma_{e}^{2})^{2}}{2\sigma_{e}^{2}}+\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)\,\mathrm{d}u$ (A.40)
$=\sigma_{e}^{2}+(\sigma_{e}^{4}+\sigma_{e}^{2})\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.41)
Therefore,
$\mathbb{E}\left[\frac{\{Y(0)-\mathbb{E}[Y(0)]\}^{2}}{1-e(X)}\right]=a_{0}^{2}\sigma_{e}^{2}+\sigma_{0}^{2}+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.42)
Adding (A.32) and (A.42), we have
$\small V=(a_{1}^{2}+a_{0}^{2})\sigma_{e}^{2}+(\sigma_{1}^{2}+\sigma_{0}^{2})+\left[a_{1}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{1}^{2}\right]\exp\left(-\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right)+\left[a_{0}^{2}\sigma_{e}^{2}(\sigma_{e}^{2}+1)+\sigma_{0}^{2}\right]\exp\left(\mu_{e}+\frac{1}{2}\sigma_{e}^{2}\right).$ (A.43)
A.7 Proof that the calculated sample size increases with $\rho^{2}$.
We rewrite $\widetilde{V}=\widetilde{V}_{\mathrm{SH}}+\rho^{2}\widetilde{V}_{\mathrm{adj}}$, where $\widetilde{V}_{\mathrm{SH}}=2+\exp(-\mu_{e}+\sigma_{e}^{2}/2)+\exp(\mu_{e}+\sigma_{e}^{2}/2)$ and
$\widetilde{V}_{\mathrm{adj}}=\frac{\sigma_{e}^{2}}{\sigma_{e\mid 1}^{2}}+\frac{\sigma_{e}^{2}}{\sigma_{e\mid 0}^{2}}-2+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 1}^{2}}-1\right]\exp\left(-\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right)+\left[\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid 0}^{2}}-1\right]\exp\left(\mu_{e}+\frac{\sigma_{e}^{2}}{2}\right).$ (A.44)
Recall that $W\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$. Then, $\sigma_{e\mid z}^{2}=\mathbb{V}[W\mid Z=z]\leq\mathbb{V}[W]=\sigma_{e}^{2}$. Hence,
$\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}\geq 1,\quad\frac{\sigma_{e}^{2}(\sigma_{e}^{2}+1)}{\sigma_{e\mid z}^{2}}-1\geq\frac{\sigma_{e}^{2}}{\sigma_{e\mid z}^{2}}-1\geq 0.$ (A.45)
Therefore, $\widetilde{V}_{\mathrm{adj}}\geq 0$.
A.8 Proof of the asymptotic variance of the general Hájek WATE estimator.
Let
$\xi_{1}^{*}=\frac{\mathbb{E}[h(X_{i};\alpha^{*})Y_{i}(1)]}{\mathbb{E}[h(X_{i};\alpha^{*})]},\quad\xi_{0}^{*}=\frac{\mathbb{E}[h(X_{i};\alpha^{*})Y_{i}(0)]}{\mathbb{E}[h(X_{i};\alpha^{*})]}$
be the expected potential outcomes given the tilting function $h(x)$ and let $\alpha^{*}$ be the true parameter value for the propensity score model $e(X;\alpha)$. Let
$\hat{\xi}_{1}=\frac{\sum_{i=1}^{N}Z_{i}Y_{i}\hat{w}_{1}(X_{i})}{\sum_{i=1}^{N}Z_{i}\hat{w}_{1}(X_{i})},\quad\hat{\xi}_{0}=\frac{\sum_{i=1}^{N}(1-Z_{i})Y_{i}\hat{w}_{0}(X_{i})}{\sum_{i=1}^{N}(1-Z_{i})\hat{w}_{0}(X_{i})},$
and
$\hat{\tau}_{w}=\hat{\xi}_{1}-\hat{\xi}_{0}.$
Then $(\hat{\xi}_{1},\hat{\xi}_{0},\hat{\alpha})$ jointly solves
$0=\frac{1}{N}\sum_{i=1}^{N}\phi(Y_{i}(1),Y_{i}(0),x_{i},Z_{i};\xi_{1},\xi_{0},\alpha)=\frac{1}{N}\sum_{i=1}^{N}\begin{pmatrix}\phi_{1}(Y_{i}(1),x_{i},Z_{i};\xi_{1},\alpha)\\ \phi_{0}(Y_{i}(0),x_{i},Z_{i};\xi_{0},\alpha)\\ \psi_{\alpha}(x_{i},Z_{i};\alpha),\end{pmatrix}$
where
$\phi=\begin{pmatrix}\phi_{1}\\ \phi_{0}\\ \psi_{\alpha}\end{pmatrix}\quad\text{with}\quad\begin{aligned} \phi_{1}&=Z_{i}\cdot w_{1}(X_{i};\alpha)\cdot\{Y_{i}(1)-\xi_{1}\},\\ \phi_{0}&=(1-Z_{i})\cdot w_{0}(X_{i};\alpha)\cdot\{Y_{i}(0)-\xi_{0}\},\\ \psi_{\alpha}&=e^{\prime}(X_{i};\alpha)\{Z_{i}-e(X_{i};\alpha)\}.\end{aligned}$ (A.46)
It is straightforward to verify that $\mathbb{E}[\phi(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})]=0$. The theory of M-estimation states that
$\sqrt{N}\begin{pmatrix}\hat{\xi}_{1}^{-}\xi_{1}^{*}\\ \hat{\xi}_{0}^{-}\xi_{0}^{*}\\ \hat{\alpha}-\alpha^{*}\end{pmatrix}\stackrel{{\scriptstyle d}}{{\rightarrow}}\mathcal{N}(0,(A^{*})^{-1}B^{*}(A^{*})^{-\top}),$ (A.47)
where $A^{*}$ and $B^{*}$ are the values of $-\mathbb{E}\left(\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\right)$ and $\mathbb{E}[\phi\phi^{\top}]$ evaluated at $(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})$ respectively.
Calculation of $A^{*}$.
By definition, we have
$\frac{\partial\phi}{\partial(\xi_{1},\xi_{0},\alpha)}\big{|}_{(\xi_{1}^{*},\xi_{0}^{*},\alpha^{*})}=\begin{pmatrix}\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}&0&\frac{\partial\phi_{1}^{*}}{\partial\alpha^{\top}}\\ 0&\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}&\frac{\partial\phi_{0}^{*}}{\partial\alpha^{\top}}\\ 0&0&\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha^{\top}},\end{pmatrix}$ (A.48)
where
$\frac{\partial\phi_{1}^{*}}{\partial\xi_{1}}$ $=-Z_{i}\cdot w_{1}(X_{i};\alpha^{*}),$
$\frac{\partial\phi_{0}^{*}}{\partial\xi_{0}}$ $=-(1-Z_{i})\cdot w_{0}(X_{i};\alpha^{*}),$
$\frac{\partial\phi_{1}^{*}}{\partial\alpha}$ $=Z_{i}\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot w_{1}^{\prime}(X_{i};\alpha^{*}),$
$\frac{\partial\phi_{0}^{*}}{\partial\alpha}$ $=(1-Z_{i})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot w_{0}^{\prime}(X_{i};\alpha^{*}),$
$\frac{\partial\psi_{\alpha}^{*}}{\partial\alpha}$ $=e^{\prime\prime}(X_{i};\alpha^{*})(Z_{i}-e(X_{i};\alpha^{*}))-\left[e^{\prime}(X_{i};\alpha^{*})\right]\left[e^{\prime}(X_{i};\alpha^{*})\right]^{\top}.$
Accordingly, we have
$A^{*}=\begin{pmatrix}a_{11}&0&a_{13}\\ 0&a_{22}&a_{23}\\ 0&0&a_{33}\end{pmatrix},$
where
$a_{11}$ $=\mathbb{E}\left[Z_{i}\cdot w_{1}(X_{i};\alpha^{*})\right]=\mathbb{E}[e(X_{i};\alpha^{*})w_{1}(X_{i};\alpha^{*})]=\mathbb{E}[h(X_{i};\alpha^{*})],$
$a_{22}$ $=\mathbb{E}\left[(1-Z_{i})\cdot{w_{0}(X_{i};\alpha^{*})}\right]=\mathbb{E}[(1-e(X_{i};\alpha^{*}))w_{0}(X_{i};\alpha^{*})]=\mathbb{E}[h(X_{i};\alpha^{*})],$
$a_{13}$ $=\mathbb{E}\left[-Z_{i}\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot w_{1}^{\prime}(X_{i};\alpha^{*})\right]=-\mathbb{E}\left[e(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot w_{1}^{\prime}(X_{i};\alpha^{*})\right],$
$a_{23}$ $=\mathbb{E}\left[-(1-Z_{i})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot w_{0}^{\prime}(X_{i};\alpha^{*})\right]=-\mathbb{E}\left[(1-e(X_{i};\alpha^{*}))\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot w_{0}^{\prime}(X_{i};\alpha^{*})\right],$
$a_{33}$ $=\mathbb{E}\left[-e^{\prime}(X_{i};\alpha^{*})e^{\prime}(X_{i};\alpha^{*})^{\top}\right].$
Calculation of $B^{*}$.
$B^{*}=\mathbb{E}[\phi^{*}\phi^{*\top}]=\begin{pmatrix}b_{11}&0&b_{13}\\ 0&b_{22}&b_{23}\\ b_{13}^{\top}&b_{23}^{\top}&b_{33}\end{pmatrix},$
where
$b_{11}$ $=\mathbb{E}\left[Z_{i}\cdot{\{Y_{i}(1)-\xi_{1}^{*}\}^{2}}\cdot{w_{1}(X_{i};\alpha^{*})^{2}}\right]=\mathbb{E}\left[{\{Y_{i}(1)-\xi_{1}^{*}\}^{2}}{e(X_{i};\alpha^{*})}w_{1}(X_{i};\alpha^{*})^{2}\right],$
$b_{22}$ $=\mathbb{E}\left[(1-Z_{i})\cdot{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\cdot{w_{0}(X_{i};\alpha^{*})^{2}}\right]=\mathbb{E}\left[{\{Y_{i}(0)-\xi_{0}^{*}\}^{2}}\{1-e(X_{i};\alpha^{*})\}{w_{0}(X_{i};\alpha^{*})}^{2}\right],$
$b_{13}$ $=\mathbb{E}\left[Z_{i}(Z_{i}-e(X_{i};\alpha^{*}))\cdot w_{1}(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right]$
$=\mathbb{E}\left[e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}\cdot w_{1}(X_{i};\alpha^{*})\cdot\{Y_{i}(1)-\xi_{1}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right],$
$b_{23}$ $=\mathbb{E}\left[(1-Z_{i})(Z_{i}-e(X_{i};\alpha))\cdot w_{0}(X_{i};\alpha^{*})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right]$
$=\mathbb{E}\left[-e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}\cdot w_{0}(X_{i};\alpha^{*})\cdot\{Y_{i}(0)-\xi_{0}^{*}\}\cdot e^{\prime}(X_{i};\alpha^{*})\right],$
$b_{33}$ $=\mathbb{E}\left[e(X_{i};\alpha^{*})\{1-e(X_{i};\alpha^{*})\}e^{\prime}(X_{i};\alpha^{*})e^{\prime}(X_{i};\alpha^{*})^{\top}\right].$
Calculation of $(A^{*})^{-1}B^{*}(A^{*})^{-\top}$.
First, we notice that
$A^{*}=\begin{pmatrix}a_{11}&0&a_{13}\\ 0&a_{11}&a_{23}\\ 0&0&a_{33}\end{pmatrix}\Longrightarrow(A^{*})^{-1}=\begin{pmatrix}a_{11}^{-1}&0&-a_{11}^{-1}a_{13}a_{33}^{-1}\\ 0&a_{11}^{-1}&-a_{11}^{-1}a_{23}a_{33}^{-1}\\ 0&0&a_{33}^{-1}\end{pmatrix}$
Therefore,
$(A^{*})^{-1}B^{*}(A^{*})^{-\top}$ $=\begin{pmatrix}a_{11}^{-1}&0&-a_{11}^{-1}a_{13}a_{33}^{-1}\\ 0&a_{11}^{-1}&-a_{11}^{-1}a_{23}a_{33}^{-1}\\ 0&0&a_{33}^{-1}\end{pmatrix}\begin{pmatrix}b_{11}&0&b_{13}\\ 0&b_{22}&b_{23}\\ b_{13}^{\top}&b_{23}^{\top}&b_{33}\end{pmatrix}\begin{pmatrix}a_{11}^{-1}&0&0\\ 0&a_{11}^{-1}&0\\ -a_{33}^{-1}c_{13}^{\top}a_{11}^{-1}&-a_{33}^{-1}c_{23}^{\top}a_{11}^{-1}&a_{33}^{-1}\end{pmatrix}$
$=a_{11}^{-2}\begin{pmatrix}b_{11}-2a_{13}a_{33}^{-1}b_{13}^{\top}+a_{13}a_{33}^{-1}a_{13}^{\top}&-a_{13}a_{33}^{-1}b_{23}^{\top}-b_{13}a_{33}^{-1}a_{23}^{\top}+a_{13}a_{33}^{-1}a_{23}^{\top}&*\\ -a_{13}a_{33}^{-1}b_{23}^{\top}-b_{13}a_{33}^{-1}a_{23}^{\top}+a_{13}a_{33}^{-1}a_{23}^{\top}&b_{22}-2a_{23}a_{33}^{-1}b_{23}^{\top}+a_{23}a_{33}^{-1}a_{23}^{\top}&*\\ *&*&*\end{pmatrix}.$
Hence,
$\mathrm{Var}(\hat{\tau}_{w})$ $=\begin{pmatrix}1&-1&0\end{pmatrix}(A^{*})^{-1}B^{*}(A^{*})^{-\top}\begin{pmatrix}1\\ -1\\ 0\end{pmatrix}$
$=a_{11}^{-2}\left[(b_{11}+b_{22}-2(a_{13}-a_{23})a_{33}^{-1}(b_{13}-b_{23})^{\top}+(a_{13}-a_{23})a_{33}^{-1}(a_{13}-a_{23})^{\top})\right].$
When the propensity scores are assumed known, $e^{\prime}(X_{i};\alpha^{*})=0$, and hence $a_{13}=a_{23}=0$. Consequently, the variance reduces to $a_{11}^{-2}(b_{11}+b_{22})$. However, unlike ATE, there is no theoretical guarantee on whether assuming known propensity score leads to a larger or a smaller variance.
A.9 Proof of (C.5)
Recall that $Y(z)=a_{z}W_{e}+\epsilon_{z}$, where $\epsilon_{z}\sim\mathsf{N}(\mu_{z},\sigma_{z}^{2})$. Then,
$\xi_{z}=\frac{\mathbb{E}[h(X)Y(z)]}{\mathbb{E}[h(X)]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\frac{\mathbb{E}[h(W_{e})\epsilon_{z}]}{\mathbb{E}[h(W_{e})]}=a_{z}\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}+\mu_{z}.$ (A.49)
The last equation holds because $\epsilon_{z}$ and $W_{e}$ are uncorrelated normal random variables and thus independent. Then,
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.50)
$=\mathbb{E}\left[\mathbb{E}\left[\left\{a_{1}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)+\epsilon_{1}-\mu_{1}\right\}^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\mid W_{e}\right]\right]$ (A.51)
$=\mathbb{E}\left[\left\{a_{1}^{2}\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}+\sigma_{1}^{2}\right\}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]$ (A.52)
$=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.53)
Similarly,
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right].$ (A.54)
B Examples
B.1 Normality of linear combination of covariates.
We consider the simulated data in Section 4 and three real world datasets: (i) RHC; (ii) the Best Apnea Interventions for Research (BestAIR) trial, a individually randomized trial with 169 units, with 9 continuous or binary covariates; more details can be found in Zeng et al., (2021); (iii) HSR: an observational study on the racial disparities of breast cancer screening, with 56,480 units and 17 binary covariates; more details can be found in Li et al., (2013). For each dataset, we fit a simple logistic model to estimate the propensity score $e(X)=\operatorname{\mathrm{e}xpit}(\beta_{0}+X^{\prime}\beta)$, and we provide the density plot and Q-Q plot in Figure 4.
Figure 4: Density plot and Q-Q plot of fitted logit propensity scores, which is a linear combination of covariates. The proximity of the points to the reference line in the Q-Q plot shows the closeness of the distribution to a normal distribution.
B.2 Examples of the Beta approximation of logit-normal distribution.
Figure 5 presents the density of the Beta distribution over a grid of $(a,b)$ values within $\{0.5,1,2,3\}$, and the density of the corresponding approximated logit-normal distributions.
Figure 5: Examples of the Beta approximation to logit-normal distribution
B.3 Calculation of overlap coefficient $\phi$ with fitted propensity scores
In real applications, researchers do not have access to data and thus cannot calculate $\phi$ with fitted propensity scores. Nevertheless, we provide a computation method here for interested readers. This method is more robust than directly estimating the density and can be useful for simulation studies.
By Bayes’ theorem,
$f_{z}(u)=\Pr(e(X)=u\mid Z=z)=\frac{\Pr(Z=z\mid e(X)=u)\Pr(e(X)=u)}{\Pr(Z=z)}.$ (B.1)
Hence,
$\phi=\int_{0}^{1}\sqrt{f_{0}(u)f_{1}(u)}\,\mathrm{d}u=\frac{\int_{0}^{1}\sqrt{u(1-u)}\Pr(e(X)=u)\,\mathrm{d}u}{\sqrt{r(1-r)}}.$ (B.2)
Let $F(u)=\Pr(e(X)\leq u)$ be the CDF of $e(X)$. Let $\widehat{e}_{1},\widehat{e}_{2},\dots,\widehat{e}_{n}$ be the estimated propensity scores of the units, and let $\widehat{F}(u)=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(\widehat{e}_{i}\leq u)$ be the empirical CDF of $e(X)$. Then,
$\phi=\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}F(u)}{\sqrt{r(1-r)}}\approx\frac{\int_{0}^{1}\sqrt{u(1-u)}\,\mathrm{d}\widehat{F}(u)}{\sqrt{r(1-r)}}=\frac{1}{n}\cdot\frac{\sum_{i=1}^{n}\sqrt{\widehat{e}_{i}(1-\widehat{e}_{i})}}{\sqrt{r(1-r)}}.$ (B.3)
C Extension to weighted average treatment effect estimands
We extend the above methodology to a general class of weighted average treatment effect (WATE), which includes the ATE and several other common estimands as special cases. Averaging the CATE $\tau(x)$ over a target population, we obtain the average treatment effect of that population. Specifically, assume the observed sample is drawn from a population with covariates distribution $f(x)$, and let $f(x)h(x)$ denote the covariate distribution of the target population, where $h(x)$ is called a tilting function. Then we can represent the average treatment effect on the target population by a WATE
$\tau^{h}=\frac{\operatorname{\mathbb{E}}[\tau(x)h(x)]}{\operatorname{\mathbb{E}}[h(x)]}.$ (C.1)
When $h(x)=1$, $\tau^{h}$ reduces to the ATE; when $h(x)=e(x)$, $\tau^{h}$ becomes the average treatment effect for the treated (ATT); when $h(x)=e(x)(1-e(x))$, $\tau^{h}$ becomes the average treatment effect for the overlap population (ATO). We can use a unified weighting strategy to identify the WATE because $\tau^{h}=\operatorname{\mathbb{E}}[w_{1}(X)ZY-w_{0}(X)(1-Z)Y]$, where $w_{1}(x)=h(x)/e(x),w_{0}(x)=h(x)/(1-e(x))$ (Li et al.,, 2018). This corresponds to (i) the inverse probability weight, $(w_{1}=1/e(x),w_{0}=1/(1-e(x)))$, for the ATE; (ii) the ATT weight, $(w_{1}=1,w_{0}=e(x)/(1-e(x)))$, for the ATT; and (iii) the overlap weight (OW), $(w_{1}=1-e(x),w_{0}=e(x))$, for the ATO.
The Hájek estimator for the WATE is:
$\widehat{\tau}_{w}=\frac{\sum_{i}w_{1}(X_{i})Z_{i}Y_{i}}{\sum_{i}w_{1}(X_{i})Z_{i}}-\frac{\sum_{i}w_{0}(X_{i})(1-Z_{i})Y_{i}}{\sum_{i}w_{0}(X_{i})(1-Z_{i})}.$ (C.2)
We can adapt the procedures in Section 3 to derive sample size calculation formulas for the ATT and ATO estimands. Specifically, the procedures to determine the propensity score and the outcome distributions remain the same, but the variance formula changes. Unlike the ATE, there is no deterministic order between the variance of the Hájek estimator with the estimated versus true propensity score for other WATE estimands (Yang and Ding,, 2018). For the same reasons as with the ATE stated in Section 2, we choose to proceed with the variance with the true $e$.
Below we derive the variance of the Hájek estimator for the WATE with the tilting function $h$:
$V_{w}\equiv\operatorname{\mathbb{V}}(\widehat{\tau}_{w})=\frac{1}{\mathbb{E}[h(X_{i})^{2}]}\mathbb{E}\left[\left\{\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}+\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}\right\}h(X_{i})^{2}\right],$ (C.3)
where $\xi_{z}=\mathbb{E}_{h}[Y_{i}(z)]=\mathbb{E}[h(X_{i})Y_{i}(z)]/\mathbb{E}[h(X_{i})]$; the proof is relegated to Appendix A.8. For the ATE, $h(X_{i})=1$, and $V_{w}$ reduces to $V$ as in Equation 3. For the ATT and ATO, $h(X_{i})$ is a function of the propensity score $e(X_{i})$ and hence $V_{w}$ is also fully determined by the joint distribution of $e(X_{i})$ and the potential outcomes $Y_{i}(z)$, which can be computed via (16). However, either the ATT or the ATO has a closed-form expression of variance, but they can be computed numerically.
We outline how to compute $V_{w}$ with the tilting function $h(X)$ given the parameters $(\mu_{e},\sigma_{e})$ and $(a_{1},a_{0},\mu_{1},\mu_{0},\sigma_{1}^{2},\sigma_{0}^{2})$. For the ATT and ATO, $h(X)$ is a function of the propensity score $e(X)$, whose distribution can be approximated by $W_{e}$. Therefore, we can also write the tilting function as $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})$ for ATT and $h(W_{e})=\operatorname{\mathrm{e}xpit}(W_{e})[1-\operatorname{\mathrm{e}xpit}(W_{e})]$ for ATO, such that $h(W_{e})$ has approximately the same distribution as $h(X)$. Calculating the denominator $\mathbb{E}[h(X)^{2}]=\mathbb{E}[h(W_{e})^{2}]$ where $W_{e}\sim\mathsf{N}(\mu_{e},\sigma_{e}^{2})$ involves numerically (e.g. Simpson’s method) evaluating the following integral
$\int_{-\infty}^{\infty}h(u)^{2}\frac{1}{\sqrt{2\pi\sigma_{e}^{2}}}\exp\left(-\frac{(u-\mu_{e})^{2}}{2\sigma_{e}^{2}}\right)\,\mathrm{d}u,$ (C.4)
which is available in most statistical software. For the numerator, we show that
$\mathbb{E}\left[\frac{(Y_{i}(1)-\xi_{1})^{2}}{e(X_{i})}h(X_{i})^{2}\right]$ $=a_{1}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{1}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.5)
$\mathbb{E}\left[\frac{(Y_{i}(0)-\xi_{0})^{2}}{1-e(X_{i})}h(X_{i})^{2}\right]$ $=a_{0}^{2}\mathbb{E}\left[\left(W_{e}-\frac{\mathbb{E}[h(W_{e})W_{e}]}{\mathbb{E}[h(W_{e})]}\right)^{2}\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right]+\sigma_{0}^{2}\mathbb{E}\left[\frac{h(W_{e})^{2}}{1-\operatorname{\mathrm{e}xpit}(W_{e})}\right],$ (C.6)
whose proof is relegated to Appendix A.9, where all the expectation terms only involve $W_{e}$ and can hence be calculated numerically by evaluating the corresponding integrals.
Similarly as the case for ATE, the result does not depend on $E_{1}$ or $E_{0}$, and only depends on $(r,\phi,\rho^{2})$ if one assumes $S_{1}^{2}=S_{0}^{2}$ and $\rho_{1}^{2}=\rho_{0}^{2}=\rho^{2}$.
D Additional Simulation and RHC Study Results
D.1 Density plot of the estimated propensity scores in the RHC study.
Figure 6: Density plot of the estimated propensity scores by treatment arm in RHC.
D.2 Summary quantities in Simulation 4.1
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00
$\phi$ 1.00 0.98 0.93 0.87 0.83 0.81
$r$ 0.50 0.50 0.50 0.50 0.50 0.50
$E_{1}$1.73 -1.90 -2.03 -2.13 -2.17 -2.19
$E_{0}$2.74 -2.56 -2.43 -2.34 -2.30 -2.27
$S_{1}^{2}$ 20.02 20.30 20.41 20.46 20.47 20.48
$S_{0}^{2}$ 20.03 19.70 19.46 19.28 19.20 19.14
$\rho_{1}^{2}$ 0.02 0.04 0.04 0.04 0.04 0.04
$\rho_{0}^{2}$ 0.03 0.03 0.02 0.02 0.02 0.02
$S^{2}$ 20.02 20.00 19.93 19.87 19.83 19.81
$\rho^{2}$ 0.02 0.03 0.03 0.02 0.02 0.02
$R^{2}$ 0.20 0.20 0.19 0.19 0.19 0.19
Table 2: Summary quantities $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, pooled summary quantities $(S^{2},\rho^{2})$ and the regression $R^{2}$ under different values of $\kappa$ in Simulation 4.1.
D.3 Simulations with binary outcomes
We repeat the simulation design as in Section 4.1, and dichotomize the potential outcome $Y_{i}(z)$ into a binary variable $Y^{b}_{i}(z)=\mathbb{I}(Y_{i}(z)>-2)$. The observed outcome $Y^{b}_{i}=Z_{i}Y^{b}_{i}(1)+(1-Z_{i})Y^{b}_{i}(0)$. The true ATE on $Y^{b}$ is $\tau^{b}=\mathbb{E}[Y^{b}_{i}(1)-Y^{b}_{i}(0)]=0.089$.
$\kappa$ 0.00 0.25 0.50 0.75 0.90 1.00
$\phi$ 1.00 0.98 0.93 0.87 0.84 0.81
$r$ 0.50 0.50 0.50 0.50 0.50 0.50
$E_{1}$ 0.53 0.51 0.50 0.49 0.48 0.48
$E_{0}$ 0.43 0.45 0.46 0.47 0.47 0.48
$S_{1}^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$S_{0}^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$\rho_{1}^{2}$ 0.00 0.03 0.03 0.03 0.02 0.02
$\rho_{0}^{2}$ 0.00 0.02 0.01 0.01 0.01 0.01
$S^{2}$ 0.25 0.25 0.25 0.25 0.25 0.25
$\rho^{2}$ 0.00 0.02 0.02 0.01 0.01 0.01
$R^{2}$ 0.13 0.13 0.12 0.12 0.12 0.12
Table 3: Summary quantities $(r,\phi,E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1},\rho_{0})$, pooled summary quantities $(S^{2},\rho^{2})$ and the regression $R^{2}$ under different values of $\kappa$ in Simulation D.3.
Assume we estimate $\tau^{b}$ using the Hájek estimator with $Y^{b}_{i}$. Dichotomizing the outcome does not affect the propensity scores, and hence $(r,\phi)$ remain the same as in that in Simulation 4.1. The remaining quantities $(E_{1},E_{0},S_{1}^{2},S_{0}^{2},\rho_{1}^{2},\rho_{0}^{2})$ and the pooled quantities $(S^{2},\rho^{2})$ are shown in Table 3 in Supplementary Material C.3. We inversely solve for the sample size to reach a nominal power of 0.80, and show the results on Table 4.
Proposed $z$-test Shook-sa & Hudgens
$\phi$ Size True PS Est PS Size True PS Est PS Size True PS Est PS
1.00 992 0.81 0.85 977 0.80 0.83 977 0.80 0.83
0.98 1019 0.80 0.85 977 0.80 0.82 1016 0.80 0.85
0.93 1166 0.80 0.85 975 0.74 0.78 1157 0.80 0.85
0.88 1528 0.83 0.85 973 0.62 0.68 1489 0.80 0.85
0.84 1972 0.83 0.87 972 0.56 0.59 1914 0.82 0.86
0.81 2460 0.86 0.89 972 0.49 0.51 2342 0.83 0.87
Table 4: Sample size calculation by the proposed method, the two-sample $z$-test and the method in (Shook-Sa and Hudgens,, 2022) to achieve nominal power of 0.80 under various degrees of overlap. The power is obtained from random sampling of the calculated size from the $N=1,000,000$ units. The Monte Carlo error for a power of 0.8 is 0.004.
The results show that the proposed method yields sample sizes that lead to powers close to 0.80, and the results are similar to Shook-Sa and Hudgens, (2022). In contrast, the two-sample $z$-test continues to lead to vastly underestimated sample size as the degree of overlap decreases.
D.4 Simulations with ATT and ATO estimands
With the same simulation design and data generating process as in Section 4.1, we focus on the ATT and ATO estimands, the true value of which are both 1. For $\kappa\in\{0,0.25,0.5,0.75,0.9,1\}$, corresponding to $\phi\in\{1.00,0.98,0.93,0.87,0.84,0.81\}$, we calculate the sample size of the Hájek estimator (C.2) for the ATT and ATO, with estimated and true propensity scores, respectively. Then we inverse the process under each scenario to calculate the power; the results are shown in Table 5. In all scenarios, the estimated sample size is close to the true value 1000. As predicted from the general theory on the WATE (Li et al.,, 2018), the ATT estimator has large variance than the ATO estimator. Consequently, given the same setting, the ATT requires larger sample size to achieve the same level of power the ATO, and the discrepancy increases as the degree of overlap decreases.
ATT Power ATO Power
$\phi$ Size True PS Est PS Size True PS Est PS
1.00 629 0.86 0.80 629 0.87 0.80
0.98 661 0.88 0.80 653 0.85 0.79
0.93 772 0.86 0.80 697 0.80 0.76
0.88 1077 0.85 0.80 739 0.75 0.76
0.84 1457 0.84 0.79 774 0.76 0.78
0.81 1900 0.82 0.79 778 0.78 0.81
Table 5: Sample size (size) calculated by the proposed method to achieve the nominal power 0.8 under $r=0.5$ and various degrees of overlap in simulations. The power of the Hájek estimators is obtained from bootstrapping 10,000 random samples of the calculated size.
As noted in Supplementary Material C, the sample size for the ATT and ATO calculated using the estimated propensity scores is not guaranteed to be conservative. However, this simulation shows that (i) the calculated sample sizes under the true and estimated propensity scores are similar, and (ii) the sample size calculated for the ATO appears to be conservative if one expects to use the estimated propensity score in the study.
D.5 RHC study with ATT and ATO estimands
The Hájek estimate of the ATT is 0.070 (standardized effect size 0.15). The sample size calculated to achieve power 0.8 is 2449. With a sample of 2449 units, the bootstrap estimate of the power with $B=1,000$ replications is 0.695 (95% CI: 0.666 to 0.724). Unlike the ATE where we overestimate the sample size, our estimate for the ATT sample size is not guaranteed conservative. The Hájek estimate of the ATO is 0.075 (standardized effect size 0.16). The sample size calculated to achieve power 0.8 is 1546. With a sample of 1546 units, the bootstrap estimate of the power with $B=1,000$ replications is 0.816 (95% CI: 0.792 to 0.840).